crispyrat:

im back with more math Rick and Nick share their food with Albert. Rick has 5 loaves of bread and Nick has 3 loaves. The three share the bread equally. Albert gives Rick and Nick 8 dollars, which they agree to share fairly. How should Rick and Nick divide the eight dollars between them?

crispyrat:

@AZ

crispyrat:

uh whos pear? im confused xp

Extrinix:

That was a plagiarized response, I believe.

TurtlePrincess:

I accidentally did that, lol my bad I was reading and accidentally clcked it

crispyrat:

its ok i'll just wait for AZ to help

AZ:

It feels so tempting to just say that Rick should get $5 and Albert get $3 but that's not the solution haha There's 8 loaves of bread and they share is equally so 8/3 = 2 2/3 and that's how much Rick, Albert and Nick each ate and how much would you have to subtract from Rick's share to make it that average and how much from Albert's share and that should be the ratio in how you divide the money between Rick and Albert

crispyrat:

what what do u mean by average xp

AZ:

oh well by average I meant how much bread they got which is 2 2/3 I said average since they shared it equally

crispyrat:

Wait srry im so confused what is ricks share then?

AZ:

so Rick and Nick come to the party with their loaves of bread and Albert comes empty handed They split the bread equally 5 + 3 = 8 loaves of bread 8 / 3 = 2.67 or 2 2/3 to be exact that's how much each one gets

AZ:

now Albert pays them 8 dollars for the bread how should they split the money? so the best way to think about it is how much did Rick give from his share to reach 2 2/3 and similarly for Nick and that's the ratio the money should be split in

crispyrat:

oh i see

AZ:

so like if I want to make an analogy let's say I have 2 cakes and then we meet up and I give you one cake so we both had one cake each you pay me $5 for the cake it's all simple since there's two people but let's say there's three people and I have 3 cakes and you brought 3 cakes and this other guy brought none we both gave him one cake each and now we all got to eat 2 cakes so if he paid us any money, we would split it equally since we both gave him one cake but let's say instead I had 4 cakes, you had 2 cakes and this third person had 0 I give 2 cakes to this person now we all get to eat 2 cakes and if he pays us, I get all the money because they were my 2 cakes

AZ:

does that make more sense

crispyrat:

so the answer is (5-8/3)*8 (3-8/3)*8

AZ:

hmm no 5 - x = 8/3 3 - y = 8/3 and then the $8 should be split in the ratio of x to y so x/(x+y) * 8 and y/(x+y) * 8

crispyrat:

5-8/3=15/3-8/3=7/3 9/3-8/3=1/3 x=7/3 y=1/3 (7/3)/(8/3)=7/8*8=7 (1/3)/(8/3)=1/8*8=1 @AZ would this work?

AZ:

ohhh so yeah, you were right about the 5 - 8/3 3 - 8/3 If I just read what I posted then I would have seen x and y is equivalent to that But that is correct. Good job!!